OpenStudy (anonymous):

If → u = 3/5 →i - 4/5→j, find D→u f (2, pi/3) DESPERATE!!

OpenStudy (anonymous):

problem 1 part b

OpenStudy (anonymous):

It want directional derivative

OpenStudy (anonymous):

\[ f(x, y) = x cos y \] let's find Gradient \[\nabla f=<cos (y),-xsin (y)>\] In the direction of \[\frac{3}{5}i-\frac{4}{5}j\] This is the unit vector since magnitude is 1 Now take dot product \[<3/5,-4/5>.<cos (y),-xsin (y)>\] \[\frac{4}{5} x \sin (y)+\frac{3 \cos (y)}{5}\] Plug in value (x,y) --> (2,pi/3) \[\frac{3}{10}+\frac{4 \sqrt{3}}{5}\]

OpenStudy (anonymous):

that's all??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh wow! Thank you!!!! have been stuck on this problem for 2 days I was thrown off the the D vector u

OpenStudy (anonymous):

u r a life saver!

OpenStudy (anonymous):

wow, you've asked this question like how many times? glad I could help